For a two tailed t-test with 20 degrees of freedom using α = 0.05, the rejection region is defined as t values below or equal to _________or t-values above or equal to ______
Answer #1

solution:

Degrees of freedom = df = 20

=0.05

  / 2= 0.05 / 2 = 0.025

t /2,df = t0.025,20 = +-2.085 ( using student t table)

t values below or equal to ___-2.085______or t-values above or equal to __2.085____